# Nonparametric Smoothing of Directional and Axial Data

### v1

@@ 3.2 Smoothing via local GLMs
with a certain weight function $w_{\boldsymbol{x}_{o}}(\cdot):\mathcal{X}\to[0,\infty)$. This leads to a target function $\boldsymbol{\Theta}\mapsto L_{\boldsymbol{x}_{o}}(\boldsymbol{\Theta})$. The only modification in the previous formulae (7) for its gradient and (8) for its Hessian matrix are additional weights $w_{\boldsymbol{x}_{o}}(\boldsymbol{X}_{\!i})$ for all summands in $\sum_{i=1}^{n}\cdots$. Often it is advisable to parametrize the family $\mathcal{F}$ or $\mathcal{F}^{o}$ locally to improve numerical stability.
Specifically let $\mathcal{X}=\mathbb{R}^{q}$. For a given number $N\in(0,n)$, one could set
@@
with the inverse $\boldsymbol{\mu}^{-1}:\{\boldsymbol{z}\in\mathbb{R}^{d}:\|\boldsymbol{z}\|<1\}\to\mathbb{R}^{d}$ of the mean function $\boldsymbol{\mu}(\cdot)$ for vMF distributions.
@@
If $\mathcal{F}$ consists of all quadratic functions $\boldsymbol{f}:\mathbb{R}^{q}\to\mathbb{R}^{d}$, then these $d+1$ local basis functions are complemented with the $q(q+1)/2$ functions
@@
Numerical example.
@@
Figure 4 depicts this function $\boldsymbol{f}^{*}$ as a vector field (left panel) and shows the raw data $(\boldsymbol{X}_{\!i},\boldsymbol{Y}_{\!\!i})$ (right panel). For the true function, at each point $\boldsymbol{x}_{o}$ on the regular grid $\mathcal{X}_{o}=\{k/10:-10\leq k\leq 10\}^{2}$ of $44121^{2}=$ 441 points (gray bullets), a line segment connecting $\boldsymbol{x}_{o}$ with $\boldsymbol{x}_{o}+0.18\cdot\boldsymbol{\mu}(\boldsymbol{f}^{*}(\boldsymbol{x}_{o}))$ is attached (black line). Similarly, at each location $\boldsymbol{X}_{\!i}$, a line segment connecting $\boldsymbol{X}_{\!i}$ and $\boldsymbol{X}_{\!i}+0.05\cdot\boldsymbol{Y}_{\!\!i}$ is attached.
@@
Figure 4: True function $\boldsymbol{f}^{*}$ (left panel) and simulated raw data $(\boldsymbol{X}_{\!i},\boldsymbol{Y}_{\!\!i})$ (right panel).
The error of a local polynomial estimator $\widehat{\boldsymbol{f}}$ was quantified by
$\mathrm{RMSE}(\boldsymbol{f})\ :=\ \sqrt{441^{-1}\sum_{\boldsymbol{x}_{o}\in\mathcal{X}_{o}}\bigl{\|}\boldsymbol{\mu}(\widehat{\boldsymbol{f}}(\boldsymbol{x}_{o}))-\boldsymbol{\mu}(\boldsymbol{f}^{*}(\boldsymbol{x}_{o}))\bigr{\|}^{2}}.$
The following table shows $\mathrm{RMSE}(\boldsymbol{f})$ (rounded to four digits) for the three types of local polynomial estimators and three different values of $N$:
$\begin{array}[]{|c||c|c|c|}\hline\cr N&\text{constant}&\text{linear}&\text{quadratic}\\ \hline\cr\hline\cr 100&0.0755&0.0875&0.1279\\ \hline\cr 200&0.0725&0.0704&0.0913\\ \hline\cr 400&0.0902&0.0753&0.0650\\ \hline\cr\end{array}$
@@
These numbers reflect only one data set, but they indicate that for $N=100$, the local constant estimator outperforms the other two. For $N=200$, the local linear estimator performs best, whereas for $N=400$, the local quadratic estimator is superior to the other two. Looking at pictures of $\widehat{\boldsymbol{f}}$ and $\boldsymbol{f}^{*}$, these effects may be explained as in local polynomial least squares regression: For smaller values of $N$, the local linear and local quadratic estimators tend to overfit the data, whereas for larger values of $N$, the bias of the local constant or local linear estimator tends to be larger than in case of local quadratic modelling. Figures 6, 7 and 8 depict the three estimators $\widehat{\boldsymbol{f}}$ for $N=400$ (black), together with the true function $\boldsymbol{f}^{*}$ (green). Looking at these plots carefully, one can see that the local constant estimator seems to be more biased than the other two close to the boundary of $[-1,1]^{2}$. On the other hand, the local constant and linear estimators tend to underestimate the norm of $\boldsymbol{\mu}(\boldsymbol{f}^{*})$ in the central region, whereas the local quadratic estimator is rather accurate there.
Figure 6: Locally constant fit $\widehat{\boldsymbol{f}}$ (black) and $\boldsymbol{f}^{*}$ (green) in case of $N=400$.
Figure 7: Locally linear fit $\widehat{\boldsymbol{f}}$ (black) and $\boldsymbol{f}^{*}$ (green) in case of $N=400$.
Figure 8: Locally quadratic fit $\widehat{\boldsymbol{f}}$ (black) and $\boldsymbol{f}^{*}$ (green) in case of $N=400$.
4 Analyzing ice cracks on Europa
@@
where $\tilde{V}\in[0,2\pi)$ is a random variable following the density (3). To relate the observations $(\boldsymbol{X}_{\!i},\boldsymbol{V}_{\!\!i})$ to the point $\boldsymbol{x}_{o}\in\mathbb{S}^{2}$ and its tangent plane $\boldsymbol{x}_{o}^{\perp}$, we need a suitable transformation which is described in the next two paragraphs.
A stereographic projection.
@@
Smoothing the data for an arbitrary reference point.
@@
Since $\boldsymbol{A}(\boldsymbol{B}_{o}\boldsymbol{X}_{\!i})\boldsymbol{B}_{o}\boldsymbol{V}_{\!\!i}$ represents an axis in $\mathbb{R}^{2}$, we replace it by means of the mapping $\boldsymbol{y}(\cdot)$ in (1) with the direction $\boldsymbol{y}(\boldsymbol{A}(\boldsymbol{B}_{o}\boldsymbol{X}_{\!i})\boldsymbol{B}_{o}\boldsymbol{V}_{\!\!i})\in\mathbb{S}^{1}$. Then we apply the local polynomial estimators described in Section 3.2 to the data pairs $\bigl{(}P(\boldsymbol{B}_{o}\boldsymbol{X}_{\!i}),\boldsymbol{y}(\boldsymbol{A}(\boldsymbol{B}_{o}\boldsymbol{X}_{\!i})\boldsymbol{B}_{o}\boldsymbol{V}_{\!\!i})\bigr{)}\in\mathbb{R}^{2}\times\mathbb{S}^{1}$ and $\boldsymbol{0}$ in place of $(\boldsymbol{X}_{\!i},\boldsymbol{Y}_{\!\!i})$ and $\boldsymbol{x}_{o}$. This yields an estimated vMF parameter $\widehat{\boldsymbol{z}}\in\mathbb{R}^{2}$. Writing $\widehat{\boldsymbol{z}}=\widehat{\kappa}[\cos(2\widehat{\beta}),\sin(2\widehat{\beta})]^{\top}$ with $\widehat{\kappa}\geq 0$ and $\widehat{\beta}\in[0,\pi)$, we replace $\widehat{\boldsymbol{z}}$ with the axis parameter $\widehat{\boldsymbol{w}}:=\widehat{\kappa}[\cos(\widehat{\beta}),\sin(\widehat{\beta})]^{\top}$ and transform it back to the Bingham parameter
@@
Some numerical results for Europa.
We analyzed the data from Europa and determined Bingham distributions $\mathrm{Bh}(\boldsymbol{x}_{o},\widehat{\boldsymbol{f}}(\boldsymbol{x}_{o}))$ for many different locations $\boldsymbol{x}_{o}$. Precisely, within each of the $19$ regions, we chose an evenly spread subset of size up to $200$ of all observed locations $\boldsymbol{X}_{\!i}$ there. For the parameter $N$ we used $100$, $200$ and $400$. Now we show some results for the particular region in Figure 2. Since the results for the three values of $N$ look very similar, we only show the ones for $N=200$ in Figure 10. The upper left panel shows only the $200$ selected points $\boldsymbol{x}_{o}$. These coincide with the selected $200$ points $\boldsymbol{X}_{\!i}$ in Figure 2, but for the computation of $\widehat{\boldsymbol{f}}$ we used all $2656$ data pairs from that region. The other three panels depict the fitted Bingham distributions by means of local constant, linear and quadratic modelling. Each distribution $\mathrm{Bh}(\boldsymbol{x}_{o},\widehat{\boldsymbol{f}}(\boldsymbol{x}_{o}))$ is represented by a blue line segment connecting the points
$\boldsymbol{x}_{o}\pm 0.1\cdot\boldsymbol{\mu}(\widehat{\boldsymbol{f}}(\boldsymbol{x}_{o}))\ =\ \boldsymbol{x}_{o}\pm 0.1\cdot\tilde{\gamma}_{2}^{\prime}(\widehat{\kappa}(\boldsymbol{x}_{o}))\widehat{\boldsymbol{u}}(\boldsymbol{x}_{o}),$
@@
Figure 10: Smoothed data from Europa with $N=200$: Locations only (top left) and fitted parameter vectors via local constant (top right), linear (bottom left) and quadratic (bottom right) modelling.
Some final comments.
@@
The simulations and data analyses were carried out with the programming language $\mathsf{R}$ (R Core Team, 2023). The code we used is availabe from the first author upon request.
5 Technical details
@@
Proof of Lemma 1.
Note that $G_{d}(t)=1+t^{2}/(2d)+t^{4}/(8d(d+2))+O(t^{6})$, whence
$\tilde{\gamma}_{d}(t)$ $=\ \log\Bigl{(}1+\frac{t^{2}}{2d}+\frac{t^{4}}{8d(d+2)}+O(t^{6})\Bigr{)}$
$=\ \frac{t^{2}}{2d}+\frac{t^{4}}{8d(d+2)}-\frac{t^{2}}{8d^{2}}+O(t^{6})$
$=\ \frac{t^{2}}{2d}-\frac{t^{4}}{4d^{2}(d+2)}+O(t^{6}),$
because $\log(1+x)=x-x^{2}/2+O(x^{3})$ as $x\to 0$. Since $\gamma(\boldsymbol{z})=\tilde{\gamma}_{d}(\|\boldsymbol{z}\|)$, this implies the asserted expansion for $\gamma(\boldsymbol{z})$. Moreover, since $\tilde{\gamma}_{d}$ is an analytic function,
$\operatorname{\mathbb{E}}(U_{t})$ $=\ \tilde{\gamma}_{d}^{\prime}(t)\ =\ \frac{t}{d}-\frac{t^{3}}{d^{2}(d+2)}+O(t^{5})\ =\ \frac{t}{d}\Bigl{(}1-\frac{t^{2}}{d(d+2)}\Bigr{)}+O(t^{5}),$
$\mathrm{Var}(U_{t})$ $=\ \tilde{\gamma}_{d}^{\prime\prime}(t)\ =\ \frac{1}{d}-\frac{3t^{2}}{d^{2}(d+2)}+O(t^{4})\ =\ \frac{1}{d}\Bigl{(}1-\frac{3t^{2}}{d(d+2)}\Bigr{)}+O(t^{4}).$
Writing $\boldsymbol{z}=t\boldsymbol{v}$ with $t=\|\boldsymbol{z}\|$ and $\boldsymbol{v}\in\mathbb{S}^{d-1}$, this shows that
$\boldsymbol{\mu}(\boldsymbol{z})\ =\ \operatorname{\mathbb{E}}(U_{t})\boldsymbol{v}\ =\ \frac{1}{d}\Bigl{(}1-\frac{\|\boldsymbol{z}\|^{2}}{d(d+2)}\Bigr{)}\boldsymbol{z}+O(\|\boldsymbol{z}\|^{5}).$
Furthermore,
$\frac{1-\operatorname{\mathbb{E}}(U_{t}^{2})}{d-1}$ $=\ \frac{1-\mathrm{Var}(U_{t})-\operatorname{\mathbb{E}}(U_{t})^{2}}{d-1}$
$=\ \frac{1}{d-1}\Bigl{(}1-\frac{1}{d}+\frac{3t^{2}}{d^{2}(d+2)}-\frac{t^{2}}{d^{2}}\Bigr{)}+O(t^{4})$
$=\ \frac{1}{d-1}\Bigl{(}\frac{d-1}{d}-\frac{(d-1)t^{2}}{d^{2}(d+2)}\Bigr{)}+O(t^{4})$
$=\ \frac{1}{d}\Bigl{(}1-\frac{t^{2}}{d(d+2)}\Bigr{)}+O(t^{4}),$
whence
$\boldsymbol{\Sigma}(\boldsymbol{z})$ $=\ \mathrm{Var}(U_{t})\,\boldsymbol{v}\boldsymbol{v}^{\top}+\frac{1-\operatorname{\mathbb{E}}(U_{t}^{2})}{d-1}\,(\boldsymbol{I}_{d}-\boldsymbol{v}\boldsymbol{v}^{\top})$
$=\ \frac{1}{d}\Bigl{(}1-\frac{3\|\boldsymbol{z}\|^{2}}{d(d+2)}\Bigr{)}\,\boldsymbol{v}\boldsymbol{v}^{\top}+\frac{1}{d}\Bigl{(}1-\frac{\|\boldsymbol{z}\|^{2}}{d(d+2)}\Bigr{)}\,(\boldsymbol{I}_{d}-\boldsymbol{v}\boldsymbol{v}^{\top})+O(\|\boldsymbol{z}\|^{4})$
$=\ \frac{1}{d}\Bigl{(}1-\frac{\|z\|^{2}}{d(d+2)}\Bigr{)}\,\boldsymbol{I}_{d}-\frac{2}{d^{2}(d+2)}\boldsymbol{z}\boldsymbol{z}^{\top}+O(\|\boldsymbol{z}\|^{4}).$
@@
$\int_{-1}^{1}(1-u)^{\ell}e^{tu}h(u)\,\mathrm{d}u\ =\ \Gamma(\ell+a)\,e^{t}t^{-(\ell+a)}\Bigl{(}\sum_{j=0}^{k}b_{j}[\ell+a+j-1]_{j}t^{-j}+O(t^{-(k+1)})\Bigr{)}$
@@
Here and throughout this section, we write $[b]_{0}:=1$ and $[b]_{j}:=\prod_{i=0}^{j-1}(b-i)$ for real numbers $b$ and integers $j\geq 1$.
@@
The asserted expansion is now a consequence of the fact that $\Gamma(\ell+a+j)=[\ell+a+j-1]_{j}\Gamma(\ell+a)$ for integers $j\geq 0$. ∎
@@
as $r\downarrow 0$. Applying Lemma 3 to $h_{d}$ with the latter expansion leads to expansions for $G_{d}(t)$ and $\log G_{d}(t)$.
@@
Then, Lemma 3 implies that
$G_{d}(t)$ $=\ \frac{2^{a_{d}-1}\Gamma(d/2)}{\Gamma(1/2)}\,e^{t}t^{-a_{d}}\Bigl{(}1-\frac{a_{d}(a_{d}-1)}{2t}+\frac{[a_{d}+1]_{2}[a_{d}-1]_{2}}{8t^{2}}+O(t^{-3})\Bigr{)}$
$=\ \frac{2^{a_{d}-1}\Gamma(d/2)}{\Gamma(1/2)}\,e^{t}t^{-a_{d}}\Bigl{(}1-\frac{[a_{d}]_{2}}{2t}+\frac{[a_{d}+1]_{4}}{8t^{2}}+O(t^{-3})\Bigr{)},$
where these and the subsequent expansions are meant as $t\to\infty$. Moreover,

### v4

@@ 3.2 Smoothing via local GLMs
with a certain weight function $w_{\boldsymbol{x}_{o}}(\cdot):\mathcal{X}\to[0,\infty)$. In principle, one could even let the model $\mathcal{F}$ depend on $\boldsymbol{x}_{o}$, but for our specific settings this is not done.
Specifically let $\mathcal{X}=\mathbb{R}^{q}$. For a given number $N\in(0,n)$, one could look for the $N$-nearest neighbors of $\boldsymbol{x}_{o}$ among the observed vectors $\boldsymbol{X}_{\!i}$ and then fit a local polynomial model to the corresponding pairs $(\boldsymbol{X}_{\!i},\boldsymbol{Y}_{\!\!i})$. That is, one could reorder the observations $(\boldsymbol{X}_{\!i},\boldsymbol{Y}_{\!\!i})$ such that the distance $\|\boldsymbol{X}_{\!i}-\boldsymbol{x}_{o}\|$ is non-decreasing in $i$, and then set $w_{\boldsymbol{x}_{o}}(\boldsymbol{x})=1_{[\|\boldsymbol{x}-\boldsymbol{x}_{o}\|\leq R_{N}]}$ with $R_{N}:=\|\boldsymbol{X}_{\!N}-\boldsymbol{x}_{o}\|$. We propose a smooth version of this nearest-neighbor approach and set
@@
with the local mean
$\bar{\boldsymbol{Y}}(\boldsymbol{x}_{o})\ :=\ \sum_{i=1}^{n}w_{\boldsymbol{x}_{o}}(\boldsymbol{X}_{i})\boldsymbol{Y}_{\!\!i}\Big{/}\sum_{i=1}^{n}w_{\boldsymbol{x}_{o}}(\boldsymbol{X}_{\!i})$
and the inverse $\boldsymbol{\mu}^{-1}:\{\boldsymbol{z}\in\mathbb{R}^{d}:\|\boldsymbol{z}\|<1\}\to\mathbb{R}^{d}$ of the mean function $\boldsymbol{\mu}(\cdot)$ for vMF distributions.
@@
If $\mathcal{F}$ consists of all quadratic functions $\boldsymbol{f}:\mathbb{R}^{q}\to\mathbb{R}^{d}$, then the former $q+1$ local basis functions are complemented with the $q(q+1)/2$ functions
@@
3.3 Numerical example and simulation study
@@
Figure 4 depicts the regression function $\boldsymbol{\mu}(\boldsymbol{f}^{*})$ as a vector field (left panel) and shows the raw data $(\boldsymbol{X}_{\!i},\boldsymbol{Y}_{\!\!i})$ (right panel). For the true function, at each point $\boldsymbol{x}_{o}$ on the regular grid $\mathcal{X}_{o}=\{k/10:-10\leq k\leq 10\}^{2}$ of $44121^{2}=$ 441 points (gray bullets), a line segment connecting $\boldsymbol{x}_{o}$ with $\boldsymbol{x}_{o}+0.18\cdot\boldsymbol{\mu}(\boldsymbol{f}^{*}(\boldsymbol{x}_{o}))$ is attached (black line). Similarly, at each location $\boldsymbol{X}_{\!i}$, a line segment connecting $\boldsymbol{X}_{\!i}$ and $\boldsymbol{X}_{\!i}+0.05\cdot\boldsymbol{Y}_{\!\!i}$ is attached.
@@
Figure 4: True function $\boldsymbol{\mu}(\boldsymbol{f}^{*})$ (left panel) and simulated raw data $(\boldsymbol{X}_{\!i},\boldsymbol{Y}_{\!\!i})$ (right panel).
@@
Figure 5 depicts estimated regression functions $\boldsymbol{\mu}(\widehat{\boldsymbol{f}})$ for $N=400$ together with the true regression function $\boldsymbol{\mu}(\boldsymbol{f}^{*})$. Precisely, one sees the estimators based on local constant and local quadratic modelling. Looking at these plots carefully, one can see that the first estimator is more biased than the second one close to the boundary of $[-1,1]^{2}$. Moreover, the first estimator seems to underestimate the norm of $\boldsymbol{\mu}(\boldsymbol{f}^{*})$ in the central region, whereas the second one is rather accurate there.
Figure 5: Estimated regression functions $\boldsymbol{\mu}(\widehat{\boldsymbol{f}})$ based on local constant (left panel) and local quadratic (right panel) modelling in case of $N=400$. In addition to the fit $\boldsymbol{\mu}(\widehat{\boldsymbol{f}})$ (black) one sees $\boldsymbol{\mu}(\boldsymbol{f}^{*})$ (green).
Next we performed a little simulation study with $100$ simulations of such a data set. With these simulations we estimated for the three types of local polynomial estimates and different values of $N$ the following quantities:
$\mathrm{BIAS}$ $:=\ \sqrt{441^{-1}\sum_{\boldsymbol{x}_{o}\in\mathcal{X}_{o}}\bigl{\|}\operatorname{\mathbb{E}}\boldsymbol{\mu}(\widehat{\boldsymbol{f}}(\boldsymbol{x}_{o}))-\boldsymbol{\mu}(\boldsymbol{f}^{*}(\boldsymbol{x}_{o}))\bigr{\|}^{2}},$
$\mathrm{SD}$ $:=\ \sqrt{441^{-1}\sum_{\boldsymbol{x}_{o}\in\mathcal{X}_{o}}\operatorname{\mathbb{E}}\bigl{[}\bigl{\|}\boldsymbol{\mu}(\widehat{\boldsymbol{f}}(\boldsymbol{x}_{o}))-\operatorname{\mathbb{E}}\boldsymbol{\mu}(\widehat{\boldsymbol{f}}(\boldsymbol{x}_{o}))\bigr{\|}^{2}\bigr{]}},$
$\mathrm{RMSE}$ $:=\ \sqrt{441^{-1}\sum_{\boldsymbol{x}_{o}\in\mathcal{X}_{o}}\operatorname{\mathbb{E}}\bigl{[}\bigl{\|}\boldsymbol{\mu}(\widehat{\boldsymbol{f}}(\boldsymbol{x}_{o}))-\boldsymbol{\mu}(\boldsymbol{f}^{*}(\boldsymbol{x}_{o}))\bigr{\|}^{2}\bigr{]}}\ =\ \sqrt{\mathrm{BIAS}^{2}+\mathrm{SD}^{2}}.$
Table 1 contains the results. These numbers show that we have similar effects as in local polynomial least squares regression. As the parameter $N$ increases, the bias increases while the variability decreases. For fixed $N$, the bias of local constant estimation is larger than the one for local linear estimation, and the latter is larger than the one for local quadratic estimation. The variability however, measured by $\mathrm{SD}$, increases with the model complexity.
[…]
Table 1: Estimated error measures for different local polynomial estimators and values $N$.
In Appendix C, we show also graphical displays of the estimated pointwise bias
$\operatorname{\mathbb{E}}\boldsymbol{\mu}(\widehat{\boldsymbol{f}}(\boldsymbol{x}_{o}))-\boldsymbol{\mu}(\boldsymbol{f}^{*}(\boldsymbol{x}_{o})).$
4 Smoothing Axial Data on a Sphere
@@
where $\tilde{V}\in[0,2\pi)$ is a random variable following the density (4). To relate the observations $(\boldsymbol{X}_{\!i},\boldsymbol{V}_{\!\!i})$ to the point $\boldsymbol{x}_{o}\in\mathbb{S}^{2}$ and its tangent plane $\boldsymbol{x}_{o}^{\perp}$, we need a suitable transformation which is described in the next two subsections.
4.1 A stereographic projection
@@
4.2 Smoothing the data for an arbitrary reference point
@@
Since $\boldsymbol{A}(\boldsymbol{B}_{o}\boldsymbol{X}_{\!i})\boldsymbol{B}_{o}\boldsymbol{V}_{\!\!i}$ represents an axis in $\mathbb{R}^{2}$, we replace it by means of the mapping $\boldsymbol{y}(\cdot)$ in (2) with the direction $\boldsymbol{y}(\boldsymbol{A}(\boldsymbol{B}_{o}\boldsymbol{X}_{\!i})\boldsymbol{B}_{o}\boldsymbol{V}_{\!\!i})\in\mathbb{S}^{1}$. Then we apply the local polynomial estimators described in Section 3 to the data pairs $\bigl{(}P(\boldsymbol{B}_{o}\boldsymbol{X}_{\!i}),\boldsymbol{y}(\boldsymbol{A}(\boldsymbol{B}_{o}\boldsymbol{X}_{\!i})\boldsymbol{B}_{o}\boldsymbol{V}_{\!\!i})\bigr{)}\in\mathbb{R}^{2}\times\mathbb{S}^{1}$ and $\boldsymbol{0}$ in place of $(\boldsymbol{X}_{\!i},\boldsymbol{Y}_{\!\!i})$ and $\boldsymbol{x}_{o}$. This yields an estimated vMF parameter $\widehat{\boldsymbol{z}}\in\mathbb{R}^{2}$. Writing $\widehat{\boldsymbol{z}}=\widehat{\kappa}[\cos(2\widehat{\beta}),\sin(2\widehat{\beta})]^{\top}$ with $\widehat{\kappa}\geq 0$ and $\widehat{\beta}\in[0,\pi)$, we replace $\widehat{\boldsymbol{z}}$ with the axis parameter $\widehat{\boldsymbol{w}}:=\widehat{\kappa}[\cos(\widehat{\beta}),\sin(\widehat{\beta})]^{\top}$ and transform it back to the Bingham parameter
@@
4.3 Some numerical results for Europa
We analyzed the data from Europa and estimated Bingham distributions $\mathrm{Bh}(\boldsymbol{x}_{o},\widehat{\boldsymbol{f}}(\boldsymbol{x}_{o}))$ at many different locations $\boldsymbol{x}_{o}$. Precisely, within each of the $19$ regions, we chose an evenly spread subset of size up to $200$ of all observed locations $\boldsymbol{X}_{\!i}$ there. Now we show some results for the particular region in Figure 2. We analyzed the data with $N=50,100,150,200,300,400$. As explained later, there is some evidence for overfitting (undersmooting) when $N=50$ and underfitting (oversmoothing) when $N=400$. For $N=100,150,200,300$, the pictures look similar and lead to the same conclusions. Figure 7 shows the estimators resulting from local linear and local quadratic models with $N=200$. For the $200$ selected points $\boldsymbol{x}_{o}$, the fitted Bingham distributions $\mathrm{Bh}(\boldsymbol{x}_{o},\widehat{\boldsymbol{f}}(\boldsymbol{x}_{o}))$ are represented by a blue line segment connecting the points
$\boldsymbol{x}_{o}\pm 0.1\cdot\tilde{\gamma}_{2}^{\prime}(\widehat{\kappa}(\boldsymbol{x}_{o}))\widehat{\boldsymbol{u}}(\boldsymbol{x}_{o}),$
@@ 4.3 Some numerical results for Europa
Figure 7: Smoothed data from Europa with $N=200$: Fitted regression function at $200$ locations via local linear (left panel) and local quadratic (right panel) models.
An interesting open problem for this type of data analysis is regression diagnostics. As a first attempt, we propose to check the plausibility of the estimators by comparing two diagnostic quantities which can be seen as a surrogate for R-squared in least squares regression: Let $\mathcal{X}_{o}=\{\boldsymbol{X}_{\!i}:i\in J_{o}\}$ be the set of $m_{o}=200$ points $\boldsymbol{x}_{o}$ in the given region at which we compute $\widehat{\boldsymbol{f}}(\boldsymbol{x}_{o})$. Then we set
$R_{\rm model}^{2}\ :=\ \frac{1}{n_{o}}\sum_{\boldsymbol{x}_{o}\in\mathcal{X}_{o}}\tilde{\gamma}_{2}^{\prime}(\widehat{\kappa}(\boldsymbol{x}_{o}))^{2}\ \in\ [0,1].$
This number measures how well the response vectors $\boldsymbol{V}_{\!\!i}$ can be predicted by the estimated regression model, assuming the latter to be true. Indeed, note that by Proposition 2,
[…]
for $\boldsymbol{V}\sim\mathrm{Bh}(\boldsymbol{x}_{o},\kappa\boldsymbol{u})$, $\kappa\geq 0$, $\boldsymbol{u}\in\mathbb{S}^{2}\cap\boldsymbol{x}_{o}^{\perp}$, where $\Psi(\boldsymbol{x}_{o},\kappa\boldsymbol{u})=\operatorname{\mathbb{E}}(\boldsymbol{V}\boldsymbol{V}^{\top})$. Thus an alternative measure of determination is given by
$R_{\rm residual}^{2}\ :=\ \frac{1}{n_{o}}\sum_{i\in J_{o}}\bigl{(}1-2\bigl{\|}\boldsymbol{V}_{\!\!i}\boldsymbol{V}_{\!\!i}^{\top}-\Psi(\boldsymbol{X}_{\!i},\widehat{\boldsymbol{f}}(\boldsymbol{X}_{\!i}))\bigr{\|}_{F}^{2}\bigr{)}.$
The ratio $R_{\rm residual}^{2}/R_{\rm model}^{2}$ should be close to one, and higher or smaller values indicate over- or underfitting, respectively.
Not surprisingly, for $N=50$, the values of $R_{\rm model}^{2}$ and $R_{\rm residual}^{2}$ are the largest, but the ratio $R_{\rm residual}^{2}/R_{\rm model}^{2}$ is $1.266$ for local constant models, which indicates overfitting. By way of contrast, for $N=400$, the same ratio is $0.952$, indicating slight underfitting. For $N=200$, the diagnostics are $(R_{\rm residual}^{2},R_{\rm model}^{2})=(0.117,0.104)$ for local linear models (ratio = 1.122) and $(R_{\rm residual}^{2},R_{\rm model}^{2})=(0.137,0.119)$ for local quadratic models (ratio = 1.152).
4.4 Final comments
@@
The simulations and data analyses were carried out with the programming language $\mathsf{R}$ (R Core Team, 2023). The code we used is availabe upon request.
Appendix
Appendix A Computational aspects of vMF distributions
The function $\gamma$.
$h_{d}(u)\ :=\ C_{d}(1-u^{2})^{a_{d}-1},\quad u\in(-1,1),$ (7)
where $a_{d}:=(d-1)/2$ and $C_{d}:=1/B(1/2,a_{d})$ with the beta function $B(\cdot,\cdot)$. This implies that $\gamma(\boldsymbol{z})=\log\int e^{\boldsymbol{z}^{\top}\boldsymbol{y}}\,M(\mathrm{d}\boldsymbol{y})$ can be written as
$\gamma(\boldsymbol{z})\ =\ \tilde{\gamma}_{d}(\|\boldsymbol{z}\|),$
where
@@ Appendix A Computational aspects of vMF distributions
Mean and Covariance of $\mathrm{vMF}(\boldsymbol{z})$.
Note first that symmetry considerations reveal that $\boldsymbol{Y}_{\!\!\boldsymbol{0}}\sim M=\mathrm{vMF}(\boldsymbol{0})$ satisfies
$\boldsymbol{\mu}(\boldsymbol{z})$ $=\ \operatorname{\mathbb{E}}(U_{t})\boldsymbol{v},$ (9)
$\boldsymbol{\Sigma}(\boldsymbol{z})$ $=\ \mathrm{Var}(U_{t})\,\boldsymbol{v}\boldsymbol{v}^{\top}+\frac{1-\operatorname{\mathbb{E}}(U_{t}^{2})}{d-1}(\boldsymbol{I}_{d}-\boldsymbol{v}\boldsymbol{v}^{\top}).$ (10)
Since the distributions of $U_{t}$, $t\in\mathbb{R}$, form an exponential family with natural parametrization, one can write $\operatorname{\mathbb{E}}(U_{t})=\tilde{\gamma}_{d}^{\prime}(t)$ and $\mathrm{Var}(U_{t})=\tilde{\gamma}_{d}^{\prime\prime}(t)$, which leads to (1). The fact that $\tilde{\gamma}_{d}^{\prime}:[0,\infty)\to[0,1)$ is bijective is well-known and follows also implicitly from Corollary 6. Note also that
with $G_{d}^{(\ell)}$ denoting the $\ell$-th derivative of $G_{d}$. Thus,
Numerical computation of $G_{d}^{(\ell)}(t)$ for moderate values of $t$.
Note first that
Consequently,
$G_{d}^{\prime}(t)\ =\ \frac{t}{d}\,G_{d+2}(t).$ (11)
From this formula, one can proceed inductively. In particular,
$G_{d}^{\prime\prime}(t)\ =\ \frac{1}{d}\,G_{d+2}(t)+\frac{t^{2}}{d(d+2)}\,G_{d+4}(t).$
@@
For large values of $t$, computing $G_{d}(t)$, $G_{d+2}(t)$ and $G_{d+4}(t)$ via their series expansion becomes problematic. Here we resort to the following approximation formulae.
Lemma 3.
@@
$\int_{-1}^{1}(1-u)^{\ell}e^{tu}h(u)\,\mathrm{d}u\ =\ e^{t}t^{-(\ell+a)}\Bigl{(}\sum_{j=0}^{k}b_{j}\Gamma(\ell+a+j)t^{-j}+O(t^{-(k+1)})\Bigr{)}$
@@
as $r\downarrow 0$. Here and throughout this section, we use the notation $[s]_{j}:=\prod_{i=0}^{j-1}(s-i)$ for real numbers $s$ and integers $j\geq 1$. Applying Lemma 4 to $h_{d}$ with the latter expansion leads to expansions for $G_{d}(t)$ and $\log G_{d}(t)$.
@@
This corollary implies the first expansion in Lemma 3.
@@
Then Lemma 4 implies that
$G_{d}(t)$ $=\ \frac{2^{a_{d}-1}\Gamma(d/2)}{\Gamma(a_{d})\Gamma(1/2)}\,e^{t}t^{-a_{d}}\Bigl{(}1-\frac{\Gamma(a_{d}+1)(a_{d}-1)}{2t}+\frac{\Gamma(a_{d}+2)[a_{d}-1]_{2}}{8t^{2}}+O(t^{-3})\Bigr{)}$
$=\ \frac{2^{a_{d}-1}\Gamma(d/2)}{\Gamma(1/2)}\,e^{t}t^{-a_{d}}\Bigl{(}1-\frac{[a_{d}]_{2}}{2t}+\frac{[a_{d}+1]_{4}}{8t^{2}}+O(t^{-3})\Bigr{)}$
where we used the identities $\Gamma(a_{d}+1)/\Gamma(a_{d})=a_{d}$ and $\Gamma(a_{d}+2)/\Gamma(a_{d})=[a_{d}+1]_{2}$. The previous and all subsequent expansions are meant as $t\to\infty$. Moreover,
@@
This corollary, applied to (9) and (10), yields the second and third expansion in Lemma 3.